Twenty-two years ago Jake was twice older than Tom. Three times Tom's age now is 16 more than twice Jake's age now. How old were they 22 years ago?
Answer by ankor@dixie-net.com(22740)

let j = Jake's present age
let t = Tom's
Write an equation for each statement
:
Twenty-two years ago Jake was twice older than Tom.
j-22 = 2(t-22)
j - 22 = 2t - 44
j = 2t - 44 + 22
j = 2t - 22
Three times Tom's age now is 16 more than twice Jake's age now.
3t = 2j+16
replace j with (2t-22)
3t = 2(2t-22) + 16
3t = 4t - 44 + 16
3t = 4t - 28
28 = 4t - 3t
t = 28 yrs is Tom's age
then
j = 2(28)- 22
j = 56 - 22
j = 34 yrs is Jake's age
:
How old were they 22 years ago?
t: 28 - 22 = 6 yrs was Tom
j: 34 - 22 = 12 yrs was Jake